OpenStudy (anonymous):

Jose put 1/2 of his money in the bank. He spent 1/3 of his remaining money on books. He has $18 left. How much did he start with? Show work, please.

OpenStudy (igreen):

Try working backwards.

OpenStudy (igreen):

\(\sf \dfrac{1}{3}\) of the money was spent on books..so divide 18 by 1/3, or multiply by its reciprocal, 3.

OpenStudy (igreen):

Tell me what you get.

OpenStudy (welshfella):

let him start with x dollars 1/2 in bank = ( 1/2) x 1/3 of the remaining = (1/3) * (1/2) * x = 1/6 x so x - ( 1/2) x - (1/6)x = 18 solve for x

OpenStudy (welshfella):

x - (2/3) x = 18 can you continue?

OpenStudy (anonymous):

IGreen, i divided 18 by 1/3, and i got 6. What do i do now?

OpenStudy (anonymous):

IGreen, i divided 18 by 1/3, and i got 6. What do i do now?

OpenStudy (welshfella):

18 / 1/3 is not 6 when diving fractions invert the divisor and multiply